OpenStudy (anonymous):

slove and show all work please. 4+square root of 4 squared-4(-1)(x) X= -------------- 2(-1)

OpenStudy (anonymous):

i think this needs an equation editor because it is not really clear

OpenStudy (anonymous):

looks like you are trying to solve something using the quadratic formula. if you post the original question it would be better.

OpenStudy (anonymous):

in particular that last x looks like it wants to be a number

OpenStudy (anonymous):

does thhis look better: -5=squared root of 4to the secound power-4 *-1(?) over 2(-1)

OpenStudy (anonymous):

\[-5=\frac{\sqrt{4^2-4\times -1\times x}}{-2}\] like this?

OpenStudy (anonymous):

yes ecpect at the bottom its 2(-1)

OpenStudy (anonymous):

i hope so. get \[10=\sqrt{16+4x}\] so \[100=16+4x\] \[84=4x\] \[x=21\]

OpenStudy (anonymous):

and befote the square root equation itss 4+

OpenStudy (anonymous):

i just wrote -2 instead of 2(-1) because they are the same. steps clear?

OpenStudy (anonymous):

yeah i get it

OpenStudy (anonymous):

good